OpenStudy (anonymous):

-5x+ 8 <= x-4 <= as in less than or equal to

OpenStudy (karatechopper):

\[-5x+ 8 \le x-4\] \[-5x+8-8\le x-4-8\] \[-5x \le x-12\] \[-5x-x \le x-x-12\] \[-6x \le -12\] \[-6x/-6 \le -12/-6\] \[x \ge 2\]